Question text Analyzing and Computing Accrued Interest on Notes Compute any interest accrued for each of the following notes payable owed by Penman, Inc., as of December 31, 2012 (assume a 365-day year). Round your answers to two decimal places. Lender Issuance Date Principal Coupon Rate (%) Term Accrued Interest Nissim 11/21/2012 $18,000 10% 120 days Answer Klein 12/13/2012 14,000 9 90 days Answer Bildersee 12/19/2012 16,000 12 60 days Answer
Answer #1

Answer :

Lender Accrued Interest
Nissim $197.26
Klein $62.14
Bildersee $63.12

Explanation:

Lender Issuance Date Principal Coupon Rate (%) Term Number of days to accrue till Dec. 31, 2012 Accrued Interest
Nissim 11/21/2012 $18,000 10% 120 days 40 days [($18,000 * 10 %* 40 days) / 365 days] = $197.26
Klein 12/13/2012 $14,000 9% 90 days 18 days [($14,000 * 9 %* 18 days) / 365 days] = $62.14
Bildersee 12/19/2012 $16,000 12% 60 days 12 days [($16,000 * 12 %* 12 days) / 365 days] = $63.12